OpenStudy (anonymous):

when f(x)=cos(x)(1+sin(x))^2 find the average rate of change of [0, pi]

OpenStudy (anonymous):

\[\frac{1}{\pi -0}\int\limits_0^{\pi} \cos(x)(1+\sin(x))^2 dx\] Then make a u substitution on 1+sin(x) and integrate.

OpenStudy (anonymous):

i got cos^3(x) does that sound right?

OpenStudy (anonymous):

It shouldn't be a function it should be a value. Let: \[\psi = 1+\sin(x) \implies d \psi = \cos(x) dx \implies \frac{1}{\pi} \int\limits_1^1 \psi^2 d \psi=0\] So this tells us the average value is zero.

OpenStudy (anonymous):

when i integrated 1+sinx i got cosx and then i plugged it back into the equation. and combined the cosx and (cosx)^2

OpenStudy (anonymous):

It doesn't work that way...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[\int\limits f(x)g(x)dx \ne \int\limits f(x)dx \int\limits g(x)dx\]

OpenStudy (anonymous):

Also \[\int\limits 1+\sin(x) dx=x-\cos(x)+C\]

OpenStudy (anonymous):

so i you the chain rule?

OpenStudy (anonymous):

? You're integrating there is no chain rule.

OpenStudy (anonymous):

I'm not seeing it. i will come back to this problem later.